# Long-Tailed 3D Detection via Multi-Modal Late-Fusion

### v4

@@ F Analysis of Hierarchical Training
  1. (a)
    Ours (e.g., Finegrain score only)
  2. (b)
    Object score * Finegrain score ([87], e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$)
  3. (c)
    Coarse score * Finegrain score (Variant-1 of [87], e.g. $P_{CAR}=P_{VEHICLE}*P_{CAR}$)
  4. (d)
    Object score * Coarse score * Finegrain score (Variant-2 of [87], e.g. $P_{CAR}=P_{OBJ}*P_{VEHICLE}*P_{CAR}$)
G Per-Class Breakdown Results
We compare recent multi-modal methods w.r.t per-class AP in Table X. All multi-modal methods perform similarly on common classes but considerably worse on Medium and Few classes, highlighting the need for further investigation of LT3D by the research community. Notably, our late-fusion approach achieves 20.0 higher AP on pushable-pullable and 6.0 higher AP on stroller than prior work. In general, our late-fusion approach yields considerable improvement on classes with medium and few examples. Despite significant improvements in rare class detection accuracy, our approach detects child with a low 8 mAP, which is still much better than compared methods. We posit that it is difficult to differentiate child from adult due to perspective geometry, since a small child close to the camera looks similar to a tall adult far away from the camera.
TABLE XII: AP vs. NDS results. We compare methods reported in the main paper w.r.t nuScenes Detection Score (NDS) and mAP. We find that the methods follow the same rankings on both metrics.
@@ G Per-Class Breakdown Results
TABLE XIII: 3D detection performance on the standard nuScenes benchmark. We evaluate the impact of late-fusion for the standard nuScenes benchmark. Although our work focuses on improving detection accuracy in-the-tail, we find that our proposed approach can still improve detection accuracy for some common classes in the standard benchmark. Notably, fusing CenterPoint [12] with DINO [11] improves detection accuracy by 2.6 mAP and 6.9 NDS. However, fusing CMT [10] (a multi-modal 3D detector) with DINO [11] increases detection accuracy by 0.2 mAP, only improving on bicycle and traffic-cone. We posit that CMT already effectively uses multi-modal training data for the (common) classes of the standard nuScenes benchmark, yielding diminishing returns with late-fusion. Note that CV stands for construction vehicle, MC represents motorcycle, and TC denotes traffic cone.
@@
H Analysis of 2D Detection Performance
Table XI shows that 2D detectors (e.g., DINO) outperform state-of-the-art 3D RGB detectors (e.g., BEVFormer) for 2D detection on the nuScenes val-set. Importantly, DINO performs significantly better than BEVFormer (15.9 vs. 2.1 mAP) on rare classes.
I Analysis of AP vs. NDS Results
We report nuScenes Detection Score (NDS) and mAP results in Table XII. We find that all prior methods follow the same ranking on both metrics. This is unsurprising because NDS is computed as a weighted sum of mAP and other true positive metrics, where mAP is weighted five times greater than other components.
J Standard nuScenes Benchmark Results
@@ H Analysis of 2D Detection Performance
We evaluate our late-fusion method on the standard 10 classes of the nuScenes benchmark [6]. Although our late-fusion approach primarily improves performance for classes with medium and few examples per class, we find that the late-fusion of CenterPoint [12] and DINO [11] yields a 2.6 and 6.9 performance increase for mAP and NDS on the standard nuScenes benchmark, respectively. Notably, we see significant perofrmance improvements for bicycle and motorcycle, likely because our LiDAR-only detector misclassifies these geometrically similar, but semantically different categories. Next We apply our late-fusion to CMT [10], a state-of-the-art multi-modal 3D detector, but find that our late-fusion approach only provides marginal performance improvements. Late-fusion of CMT [10] and DINO [11] only marginally improves bicycle and traffic cone performance. This suggests that CMT is already able to effectively utilize multi-modal information to correctly localize and classify common classes.
K State-of-the-Art Comparison on AV2
We present results on the large-scale Argoverse 2 (AV2) dataset developed for autonomous vehicle research. AV2 evaluates on 26 classes, which follow a long-tailed distribution. Following [9], we train on and evaluate detections up to 50m. As shown in Table XIV, our main conclusions from nuScenes still hold for AV2. FCOS3D [13] yields poor performance on all classes, likely due to inaccurate depth estimates. CenterPoint performs considerably better, achieving high accuracy on classes with many examples. Notably, CenterPoint performs better on AV2’s rare classes (30.2 mAP) compared to nuScenes’s rare classes (3.5 mAP), likely because AV2 has more examples per-class in-the-tail. Lastly, our proposed late-fusion approach yields a 8.3 mAPimprovement over CenterPoint and a 3.9 mAP improvement over prior work. These new results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
Few       Medium      Many
@@
TABLE XIV: Comparison with the Argoverse 2 state-of-the-art. We present results on AV2 evaluated at 50m. FCOS3D [13] achieves poor performance, likely due to inaccurate depth estimates. In contrast, CenterPoint achieves strong performance on all classes. Our multi-modal fusion approach significantly improves over CenterPoint, achieving an 8.3 mAP improvement averaged over all classes. These results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
Method Modality All Many Medium Few
FCOS3D [13] C 14.6 27.4 17.0 7.8
CenterPoint [12] L 44.0 77.4 46.9 30.2
MMF(CenterPoint, FCOS3D) (Ours) C + L 48.4 79.0 51.4 35.3
MMLF(CenterPoint, DINO) (Ours) C + L 52.3 89.4 54.2 38.7

### v5

@@ Appendix I Analysis of 2D Detection Performance
Table 13: Evaluating 2D detection performance. We find that simply projecting 3D detections from RGB-only 3D detectors to 2D image plane yields considerably lower 2D detection mAP across all classes. However, 2D RGB detectors achieve higher performance. Intuitively, 2D RGB detectors achieve better classification performance and predict tighter fitting bounding boxes than those derived by projecting 3D detections to the image plane. This explains why fusing detections from 2D RGB detectors yields better LT3D performance than detections from 3D RGB detectors. All detectors in this table are only trained on the nuScenes train-set.
Table 13 shows that using the 2D detector DINO outperforms state-of-the-art 3D RGB detectors (e.g., BEVFormer) for 2D detection on the nuScenes val-set. Importantly, DINO performs significantly better than BEVFormer (15.9 vs. 2.1 mAP) on rare classes. Furthermore, Table 11 compares using DINO [88] vs. YOLOv7 [70] as the 2D detector, applied along with the 3D LiDAR detector CenterPoint [82] in our MMLF. Results show that using DINO, a stronger detector than YOLOv7, yields significantly better 3D detection performance on all classes, particularly on medium and few classes.
Appendix J Analysis of AP vs. NDS Results
We report nuScenes Detection Score (NDS) and mAP results in Table 14. We find that all prior methods follow the same ranking on both metrics. This is unsurprising because NDS is computed as a weighted sum of mAP and other true positive metrics, where mAP is weighted five times greater than other components.
@@
Table 14: AP vs. NDS results. We compare methods reported in the main paper w.r.t nuScenes Detection Score (NDS) and mAP. We find that the methods follow the same rankings on both metrics.
Appendix K Standard nuScenes Benchmark Results
@@
Table 15: 3D detection performance on the standard nuScenes benchmark. We evaluate the impact of late-fusion for the standard nuScenes benchmark. Although our work focuses on improving detection accuracy in-the-tail, we find that our proposed approach can still improve detection accuracy for some common classes in the standard benchmark. Notably, fusing CenterPoint [82] with DINO [88] improves detection accuracy by 2.6 mAP and 6.9 NDS. However, fusing CMT [80] (a multi-modal 3D detector) with DINO [88] increases detection accuracy by 0.2 mAP, only improving on bicycle and traffic-cone. We posit that CMT already effectively uses multi-modal training data for the (common) classes of the standard nuScenes benchmark, yielding diminishing returns with late-fusion. Note that CV stands for construction vehicle, MC represents motorcycle, and TC denotes traffic cone.
@@
In Table 15, we evaluate our late-fusion method on the standard 10 classes of the nuScenes benchmark [5]. Although our late-fusion approach primarily improves performance for classes with medium and few examples per class, we find that the late-fusion of CenterPoint [82] and DINO [88] yields a 2.6 and 6.9 performance increase for mAP and NDS on the standard nuScenes benchmark, respectively. Notably, we see significant perofrmance improvements for bicycle and motorcycle, likely because our LiDAR-only detector misclassifies these geometrically similar, but semantically different categories. Next We apply our late-fusion to CMT [80], a state-of-the-art multi-modal 3D detector, but find that our late-fusion approach only provides marginal performance improvements. Late-fusion of CMT [80] and DINO [88] only marginally improves bicycle and traffic cone performance. This suggests that CMT is already able to effectively utilize multi-modal information.
Appendix L Impact of Occlusion and Range
We report average recall at four visibility levels in Table 16. Results show that detecting occluded and partially occluded objects is significantly more challenging than detecting un-occluded objects, likely because the visible regions of partially occluded objects can be easily mistaken as part of the background and hence lead to more false positive detections. Interestingly, the LiDAR-based 2D detector CenterPoint achieves the highest recall across all occlusion levels.
We also examine detection accuracy across three distance ranges in Table 17 and observe that detecting objects closer to the ego-vehicle is easier due to the higher LiDAR point density. While CenterPoint’s performance declines significantly beyond 10m, CMT and MMLF demonstrate greater robustness, suggesting that both feature-fusion and late-fusion effectively mitigate the impact of LiDAR sparsity at range. Furthermore, our multi-modal fusion approach significantly improves CenterPoint’s detection accuracy in the 10 – 20m and 20 – 30m ranges by reducing false positives, increasing AP from 47.9 to 63.2 and from 36.0 to 44.5, respectively.
Method Modality 0-40% 40-60% 60-80% 80-100%
CenterPoint (CP) L 66.0 78.7 79.5 80.1
CMT C + L 56.5 73.5 75.0 75.8
MMLF(CP, DINO) C + L 49.1 67.3 68.3 78.7
Table 16: Average recall for occluded targets at different visibility levels. We evaluate average recall at four visibility levels (defined by nuScenes). Perhaps unsurprisingly, detection accuracy drops as objects become more occluded. Note that visibility attributes are not predicted by our model and are manually annotated for the ground truth, making it difficult to measure precision.
Method Modality 0-10m 10-20m 20-30m Full Range
CenterPoint (CP) L 60.3 47.9 36.0 40.4
CMT C + L 56.9 51.5 39.8 44.4
MMLF(CP, DINO) C + L 63.2 63.2 44.5 51.4
Table 17: Detection accuracy at different distance thresholds. We evaluate detection accuracy at three distance ranges. Detecting objects closer to the ego-vehicle is easier (due to higher LiDAR point density) than detecting objects further away.
Appendix M Analysis of Recall
As shown in Table 18, the LiDAR-only 3D detector CenterPoint achieves higher recall than CMT (multi-modal) for the Pedestrian (LCA=1), Movable (LCA=1), and Object (LCA=2) superclasses. The lower recall of CMT can be attributed to its limited number of object queries (N=300) per frame compared to the (nearly) unbounded number of predictions from anchor-free detectors like CenterPoint. Importantly, the high recall of CenterPoint is essential for our late-fusion approaches (cf. Fig. 6).
Method Modality Vehicle Pedestrian Movable Object
CenterPoint L 85.5 95.5 90.7 89.3
CMT C + L 86.4 91.2 87.2 87.9
Table 18: Average recall between detectors. CenterPoint (LiDAR-only) achieves higher recall than CMT for the Pedestrian (LCA=1), Movable (LCA=1), and Object (LCA=2) superclasses.
Appendix N Impact of Calibration Noise
Translation Noise $N_{T}$ 0.01m 0.02m 0.03m 0.04m
CMT 44.4 44.4 44.4 44.4
MMLF(CP, DINO) 51.4 51.5 51.5 51.5
Rotation Noise $N_{R}$ 1 2 3 4
CMT 44.3 43.1 42.7 41.1
MMLF(CP, DINO) 48.7 47.8 46.1 45.6
Translation Noise $N_{T}$ 0.01m 0.02m 0.03m 0.04m
Rotation Noise $N_{R}$ 1 2 3 4
CMT 44.3 43.8 43.1 42.2
MMLF(CP, DINO) 47.8 46.6 45.1 44.6
Table 19: Impact of calibration noise. We apply a small amount of additive Gaussian noise to the translation and rotation calibration matrices to measure the impact of imprecise calibration on late fusion. While translation noise has little effect on detection accuracy, rotation noise significantly impacts performance. As expected, combining both translation and rotation noise degrades accuracy further.
Following prior literature, we benchmark on datasets (nuScenes and Argoverse 2) collected in-the-wild, which already contain realistic calibration errors. However, we study the impact of introducing more aggressive calibration errors in Table 19 by applying synthetic perturbations to the camera-to-LiDAR translation and rotation matrices. We apply additive Gaussian noise $N_{T}$ to each of the $X$, $Y$ and $Z$ components of the translation vector and apply additive noise $N_{R}$ to the yaw, pitch and roll of the rotation matrix. We find that translation noise does not significantly affect detection accuracy. In contrast, rotation noise has a much larger impact on accuracy. Unsurprisingly, combining translation and rotation noise yields significantly degraded accuracy. Such calibration errors affect all (early, mid, and late) fusion methods since they require calibrated extrinsics [47]. We find that our late-fusion approach is less robust to noise than end-to-end fusion methods. However, our approach still achieves higher accuracy than CMT.
Appendix O 3D Visualization of Select Figures
In Fig. 14 and 15, we visualize the LT3D detection results in a 3D virtual scene. The results are also visualized in the main paper (corresponding to Fig. 9 and 10, respectively) in the BEV scene.
MMLF(CenterPoint, DINO)           CenterPoint
@@
Figure 14: Three examples demonstrate how our multi-modal late-fusion (MMLF) approach improves LT3D by ensembling 2D RGB detections (from DINO [88]) and 3D LiDAR detections (from CenterPoint [82]). In all examples, MMLF correctly relabels detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child. This figure complements Fig. 9 in the main paper.
MMLF(CenterPoint, DINO)           TransFusion
Figure 15: Both our MMLF method (column 1) and TransFusion [2] (column 2) share the same failure cases. In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them. As a result, our MMLF misses these cars because it throws away unmatched RGB detections which do not have accurate 3D information. In the third row, we see that, although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker. As a result, the final detection is incorrect w.r.t the predicted label. TransFusion also misclassifies this object as an adult. This figure complements Fig. 10 in the main paper.